OpenStudy (thecuteone25):

Please help will MEDAL anyone who helps me the most will FAN who ever come on my question

OpenStudy (thecuteone25):

What is the area of the trapezoid? A. 44 square inches B. 168 square inches C. 64 square inches D. 192 square inches

OpenStudy (anonymous):

i can help

OpenStudy (thecuteone25):

I think that it's B

OpenStudy (study312):

which one would you like to help?

OpenStudy (thecuteone25):

anybody can help as long as I get the right answer

OpenStudy (anonymous):

192

OpenStudy (anonymous):

d

OpenStudy (thecuteone25):

can you explain to me how you got that plz

OpenStudy (study312):

one sec

OpenStudy (study312):

we want to cut so we get two triangles and 1 rectangle

OpenStudy (anonymous):

area = base1 + base2 divide ny 2 muliply by height

OpenStudy (study312):

so 18 - 14 = the base for the triangle

OpenStudy (anonymous):

area = base1 + base2 divide ny 2 muliply by height

OpenStudy (study312):

18 - 14

OpenStudy (thecuteone25):

4

OpenStudy (study312):

4 x 2 =

OpenStudy (study312):

i mean 4 x 12

OpenStudy (thecuteone25):

8

OpenStudy (anonymous):

A=a+b _______ x 2 2

OpenStudy (anonymous):

D

OpenStudy (thecuteone25):

but which bass @TheCleverOne

OpenStudy (study312):

4 x 12 = 48 and now we divide by 2, 48 ÷ 2 - 24 now the other triangle is 24

OpenStudy (mathmale):

The easiest way by far to do this problem is to find the AVERAGE length of the figure. Two parallel sides are given; they are 14 and 18 inches. What is the average of that? Next, what is the width of the figure?? The area is (average length) * (width). Please show your work..

OpenStudy (mathmale):

Hint: the average of 3 and 4 is (3+4)/2 = 3.5

OpenStudy (study312):

A=a+b 2h=14+18 2·12=192

OpenStudy (study312):

- google

OpenStudy (thecuteone25):

oh ok I get thanks guys I think I got it from here

OpenStudy (study312):

So D is right

OpenStudy (anonymous):

that is what i said

OpenStudy (thecuteone25):

oh ok thanks for helping @Study312 and @TheCleverOne

OpenStudy (anonymous):

Care to fan and medal?:)

OpenStudy (thecuteone25):

yup

OpenStudy (anonymous):

Thanks!!:)

OpenStudy (thecuteone25):

np :)